The purchase of a new gym requires an investment of 650,000 in new equipment, which will be depreciated straight-line to a zero book value over its three-year life. The market value of the equipment at the end of the project is expected to be 30% of its original cost. Annual sales from this project are estimated at $225,000 with cash expenses (not including depreciation) of $130,000. Net working capital level in each year is predicted to be 12% of the following year’s sales. The tax rate is 25%.

What is the total free cash flow in Year 0 and what is the operating cash flow in Year 1
Answer #1

Year 0:

There is negative cash flow (means outflow), since there is an aggregate of investment in equipment and net working capital for year 1 (the following year).

Cash flow = Investment + Net working capital

                 = - [650,000 + (225,000 × 12%)]

                 = - [650,000 + 27,000]

                 = - $677,000 (Answer)

Year 1:

Depreciation expense = (Investment cost – (Investment cost × 30%)) / Life years

                                    = (650,000 – (650,000 × 30%)) / 3

                                    = (650,000 – 195,000) / 3

                                    = 455,000 / 3

                                    = 151,666

Taxable income = Annual sales – (Cash expenses + Depreciation)

                           = 225,000 – (130,000 + 151,666)

                           = 225,000 – 281,666

                           = - 56,666

Tax = Taxable income × 25%

       = - 56,666 × 25%

       = - 14,167

Cash flow = - Net working capital + (Taxable income – Tax + Depreciation)

                        = - 27,000 + (- 56,666 – (-14,167) + 151,666)

                        = - 27,000 + (- 56,666 + 14,167 + 151,666)

                        = - 27,000 + 109,167

                        = $82,167 (Answer)